OpenStudy (anonymous):

What is the sum of a 22–term arithmetic sequence where the first term is 56 and the last term is –28?

OpenStudy (anonymous):

\[ a_1 =56\\ a_2 = 56 -r\\ a_3= 56 -2r\\ a_{22} = 56- 21\ r=-28\\ 21 r = 56 + 28=84\\ r= 4 \]

OpenStudy (anonymous):

\[ sum = 56 -r + 56-2r +\cdots + 56 - 21 r\\ 21(56) -4(1+2 +\cdots +21)=1176-4 (21)(22)/2=252 \]